OpenStudy (anonymous):

Catelynn is using a computer program to graph the function T(w) = –39w – 915. She wants to see the intercepts. Which minimum values should she use to set up the viewing window? (Points : 1) X Min: –1000, Y Min: –30 X Min: –900, Y Min: –50 X Min: –50, Y Min: –900 X Min: –30, Y Min: –1000

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

find the y intercept 0=-39x-915 915=-39x x=-23.46 So minimum X should be -30 to see the y-intercept. The x-intercept T(w)=-39(0)-915 T(w)=-915 So out of the options, the number that is larger that -915, is -1000. So it is A, X Min: –1000, Y Min: –30 with X Min: –1000 we will see the y-intercept -915 and with Y Min: –30 we will see the x-intercept -23.46

OpenStudy (solomonzelman):

please refresh. I am sorry about this.

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

@SolomonZelman better now?

OpenStudy (anonymous):

@Kit_Kat_Nat_<3